OpenStudy (anonymous):

Write the equation of a line in point slope form that has a slope of -3/4 and passes through the point (4,5)

OpenStudy (hba):

Use the formula \[y-y_1=m(x-x_1)\] Where, \[m=-3/4,(x_1,y_1)=(4,5)\]

OpenStudy (hba):

@undeadboss got it dude ?

OpenStudy (anonymous):

yeah im trying

OpenStudy (anonymous):

so the slope is y? right?

OpenStudy (anonymous):

m represents the slope, in this ques m=-3/4

OpenStudy (anonymous):

i had a brain fart and now i cant figure it out

OpenStudy (anonymous):

whats the equation?

OpenStudy (anonymous):

Alternate method, Equation for a line in slope form is given by, y=mx +c where m represents slope and c represents a constant. Let the line y=mx+c pass from given points (4,5) So now we have x=4 and y=5, also m=>slope=-3/4 now put these values in the equation of line as they must satisfy this eqn, from this, you will get the value of C now put the value of c and m in y=mx+c, you will get the required equation.

OpenStudy (anonymous):

i mean the end result

OpenStudy (anonymous):

try this method, u will get the ans

OpenStudy (anonymous):

the former method by hba is derived from this only.

OpenStudy (anonymous):

y-b = m(x-a) is this the right equation

OpenStudy (anonymous):

this also right, same as one by hba here a,b represent the point passing through the line and m represents the slope so, a=4, b=5 and m=-3/4 Put these value and you will get the desired eqn

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yep thx